# Anchored Preference Optimization and Contrastive Revisions: Addressing Underspecification in Alignment

### v4

@@ 5.3.1 Preference Data
The LC-AlpacaEval2.0 results generally follow a similar trend compared to MixEval-Hard, although the on-policy judge dataset attains a higher score compared to CLAIR. While both benchmarks correlate highly with human ratings of models, MixEval-Hard is our primary and most significant evaluation tool – we are able to evaluate every model checkpoint across two MixEval-Hard splits due to its low cost. Additionally, we remark on a potential issue with the robustness of LC-AlpacaEval2.0 in Appendix C. A performance breakdown in function of MixEval-Hard’s constituent benchmarks is given in Appendix B.
@@ Appendix B MixEval-Hard Performance Breakdown
Appendix C How well does AlpacaEval control for response lengths?

### v5

@@ 5.3.1 Preference Data
The LC-AlpacaEval2.0 results generally follow a similar trend compared to MixEval-Hard, although the on-policy judge dataset attains a higher score compared to CLAIR. While both benchmarks correlate highly with human ratings of models, MixEval-Hard is our primary and most significant evaluation tool – we are able to evaluate every model checkpoint across two MixEval-Hard splits due to its low cost. Additionally, we remark on a potential issue with the robustness of LC-AlpacaEval2.0 in Appendix D. A performance breakdown in function of MixEval-Hard’s constituent benchmarks is given in Appendix B.
@@ Appendix B MixEval-Hard Performance Breakdown
Appendix C Unpaired APO
In this work, we designed datasets and alignment objectives for paired preferences (output $y_{l}\prec y_{w}$ for input $x$). The original KTO objective (Ethayarajh et al., 2024) was designed to operate on desirability data (output $y$ for input $x$ was desirable or not), which does not use such paired preferences. We consider an unpaired variant of our APO-zero loss, called APO-zero-unpaired, which resembles the KTO objective but which fixes the KL term to zero. Table 5 compares KTO with APO-zero-unpaired, keeping everything else comparable with our main results in Table 2. To turn our paired datasets into unpaired datasets, we turn each datapoint consisting of two outputs into two datapoints with one output.
There is no clear winner between KTO and APO-zero-unpaired across the board. Within each dataset however, there always is a clear winner. This reflects the main findings of our work, different alignment objectives have distinct semantics, and different datasets require different semantics. APO-zero-unpaired consistently trains faster, due to not calculating the KL term. In some cases, the KTO objective can take 60% longer to train.
ME-Hard 2024-06-01 ME-Hard 2024-08-11
Dataset Objective Max $\Delta$ Mean $\Delta$ Max $\Delta$ Mean $\Delta$ Train Time
Judge KTO 2.10 --2.70 (1.67) 4.75 1.31 (1.61) 19h 18m 10s
off-policy APO-zero-unpaired --0.40 --3.67 (1.68) 4.35 0.66 (1.44) 12h 32m 58s
Judge KTO 3.50 1.28 (1.11) 4.85 2.70 (1.35) 19h 40m 10s
on-policy APO-zero-unpaired 4.35 1.31 (1.44) 5.60 3.92 (0.99) 13h 49m 55s
CLAIR KTO 3.75 1.47 (1.39) 5.80 4.12 (1.09) 17h 33m 24s
APO-zero-unpaired 1.40 --1.49 (1.77) 3.20 1.13 (1.21) 12h 31m 03s
Stronger KTO --3.25 --4.73 (1.01) 0.30 --1.18 (0.75) 19h 07m 29s
Preferred APO-zero-unpaired --2.70 --4.57 (1.32) 2.95 0.50 (1.25) 12h 38m 49s
Table 5: Max and mean MixEval-Hard improvements for the 2024-06-01 and 2024-08-11 splits, aggregated over 18 epochs of aligning Llama-3-8B-Instruct. Best overall performance bold, best performance per dataset underlined, standard deviation in parentheses. KTO is the best unpaired loss given the off-policy Judge and CLAIR datasets, while APO-zero-unpaired performs better when given the on-policy Judge and Stronger Preferred datasets. KTO can take 60% longer to train for the same configuration.
Appendix D How well does AlpacaEval control for response lengths?